# FLEX: Continuous Agent Evolution via Forward Learning from Experience

### v1

@@ 1 Introduction
Autonomous agents have achieved remarkable advances across diverse domains, including code generation (e.g., Claude Code [anthropic2025claude]), scientific discovery (e.g., Biomni [huang2025biomni], STELLA [jin2025stellaselfevolvingllmagent], and SciMaster [chai2025scimastergeneralpurposescientificai]), and in-depth research (e.g., Tongyi, OpenAI, and Gemini DeepResearch [tongyidr, openai2025deepresearch, google2025gemini]), showcasing their immense potential to solve complex, open-ended real-world tasks end to end.
However, pretrained agents remain static with frozen parameters, fundamentally precludes learning from on-the-fly experience concluded during trial-and-error, causing substantial performance degradation on challenging or untrained tasks [wu2023autogen, liu2023agentbench]. While gradient-based learning has underpinned model optimization [rumelhart1986learning, ruder2016overview, hu2022lora], it is inherently ill-suited for continuous agent evolution due to three key obstacles: (i) back-propagation incurs prohibitive computational costs [goodfellow2013empirical, cai2025training]; (ii) parameter-centric paradigm suffers from catastrophic forgetting [abbas2023loss, hadi2023large], failing to effectively learn form accumulated experiences over times; (iii) the closed-source nature of most state-of-the-art LLMs [google2025gemini2_5_pro, openai2025gpt5, anthropic2025sonnet4.5, anthropic2025claude4, xai2025grok4] renders direct parameter optimization infeasible.
Recent self-evolving paradigms attempt to overcome this limitation by evolving non-parametric components (i.e., prompts [yuksekgonul2024textgrad, zhang2025agentic], workflows [lin2025se, zhang2025multi, shang2024agentsquare], and tools [qian2023creator, qiu2025alita]) through environmental feedback and textual gradient [yuksekgonul2024textgrad, shinn2023reflexion, zhang2024revolve]. Nevertheless, these approaches encounter three vital bottlenecks towards continuous agent evolution: (i) their evolving components are task-specific, thus they cannot generalize across tasks, limiting cross-task evolution. (ii) Their evolving components (i.e. prompts, workflows, and tools) are not continuously scalable, thus they can only leverage a limited set of experiences, preventing performance from scaling with accumulated knowledge. (iii) Their evolving components are model-specific, thus new agents must interact from scratch during deployment, unable to continuously evolve from agents’ previous off-policy experiences, leading to substantial computational overhead.
@@
To validate the general effectiveness of FLEX paradigm, we conduct extensive experiments across diverse challenging scientific domains, including Olympiad-level mathematics (AIME25 [aime2025problems]), chemical retrosynthesis (USPTO50k [schneider2016s]), and protein fitness prediction (ProteinGym [notin2023proteingym]). FLEX demonstrates substantial and consistent improvements on these tasks, from $40\%$ to $63\%$ on AIME25, $20\%$ to $30\%$ on USPTO50k, and $46\%$ to $60\%$ on ProteinGym, exhibiting great enhancement in the capacity of reasoning and knowledge leverage.
@@ 3.1 Extensive Experience Exploration
To enable extensive high-quality exploration, we design a dual-scaling mechanism that combines parallel and sequential scaling. For parallel scaling, multiple reasoning trajectories are sampled in parallel for the same input through rejection sampling [fujimoto2021minimalist, dong2023raft, gulcehre2023reinforced], allowing diverse reasoning hypotheses to be explored and high-quality outputs to be retained. For sequential scaling, the actor iteratively improves upon its previous output under the critic’s semantic feedback [yuksekgonul2024textgrad, shinn2023reflexion], where the critic provides constructive guidance and refinement cues derived from the prior round. This cooperative exploration strategy balances breadth and depth, ensuring both coverage of alternative reasoning paths and progressive improvement over iterations.
@@
To ensure that collected experiences contribute meaningfully to learning, we design an iterative refinement mechanism driven by semantic evaluation. For reasoning trajectories that lead to incorrect conclusions, the critic first identifies the underlying causes of failure and summarizes them into abstract, task-agnostic improvement suggestions. These generalized feedback signals, denoted as new experiences $E_{t}$, are fed back into the actor together with its previous reasoning trace to initiate the next actor–critic iteration. This process continues until the actor produces a correct or sufficiently improved reasoning trajectory, or a predefined iteration limit is reached. In this loop, the critic’s feedback functions as a semantic update signal [yuksekgonul2024textgrad, shinn2023reflexion, madaan2023self] that guides the actor’s reasoning refinement, analogous to the gradient signal in traditional optimization, yet expressed in natural language semantics rather than numerical differentials. Unlike backpropagation-based learning, which relies on scalar gradients with limited interpretive bandwidth and heavy computational budget, this forward semantic feedback mechanism provides rich interpretable guidance without costly parameter updating, aligning with the cognitive characteristics of human-like learning from experience.
@@ 3.2 Experience Library Evolution
During inference, the experience library is encapsulated as an interactive tool callable by the agent [yao2022react, qiu2025alita]. Instead of static vector-based semantic-similarity search [reimers2019sentence, ram2023context, karpukhin2020dense], the agent performs contextualized retrieval conditioned on the specific query and its current reasoning state. This allows the agent to interpret relevance not only by semantic similarity but also by contextual consistency and goal alignment, thereby avoiding the pitfalls of retrieving semantically similar yet pragmatically contradictory entries. Retrieval proceeds hierarchically, the agent first identifies relevant high-level strategies, then locates corresponding procedural patterns, and finally accesses concrete case examples within each pattern. At each stage, the top-$k$ (typically $k=5$) most relevant entries are selected to balance precision and efficiency. The retrieved experiences support the agent’s ongoing reasoning, reflective self-evaluation, and corrective refinement. Moreover, retrieval is not restricted to pre-inference access, instead, the agent can dynamically query the library during reasoning, enabling adaptive integration of prior knowledge into evolving cognitive trajectories.
@@ 4.1 Experimental Setup
We evaluate the proposed FLEX paradigm across four challenging benchmarks spanning three scientific domains (i.e., mathematics, chemistry, and biology) to assess both reasoning capability and domain adaptability. Specifically: (1) AIME25 competition dataset [aime2025problems] for complicated mathematical reasoning; (2) GSM8k dataset [cobbe2021gsm8k] multi-step arithmetic reasoning; (3) USPTO50k [schneider2016s] benchmark for canonical single-step retrosynthesis; (4) ProteinGym [notin2023proteingym] benchmark for protein fitness prediction.
@@
Our evaluation spans a range of base models, tested on four distinct experimental configurations for comparison: (1) Vanilla LLM, (2) LLM with In-Context Learning (ICL) [dong2022survey], (3) LLM agent with reasoning-and-acting workflow (Agent) [yao2023reactsynergizingreasoningacting], and (4) our proposed FLEX. For FLEX, each model constructs an experience library during training, which is later utilized at test-time. The library structure is adapted to the benchmark: hierarchical for AIME25 and USPTO50k, and non-hierarchical for GSM8k and ProteinGym.
@@
For AIME25, the training was conducted on 49 historical problems from AIME83 to AIME24. On GSM8k, we adhere to the official training and testing splits [cobbe2021gsm8k]. For USPTO50k, 100 instances are randomly sampled from the original 5k test split for evaluation, and 50 instances from the training split are randomly sampled for training. For ProteinGym, 100 mutational sequences are sampled for each wild-type protein target as training set, which is on average 1.47% of the available data. This design reflects practical constraints in protein engineering, where only a limited number of labeled mutations are known, while the vast majority of sequence space remains unexplored.
@@ 4.2 Main Results
AIME25 Claude-Sonnet-4 [anthropic2025claude4] 40.0 30.0 (-10.0) 50.0 (+10.0) 63.3 (+23.3)
DeepSeek-V3.1-Terminus [DeepSeekV3.1Terminus] 56.7 53.3 (-3.3) 60.0 (+3.3) 66.6 (+10.0)
GSM8k GPT-3.5 [OpenAI_GPT3.5_Turbo_docs] 80.8 81.4 (+0.6) 78.5 (-2.3) 83.3 (+3.3)
GPT-4 [openai2024gpt4technicalreport] 93.8 94.2 (+0.4) 94.0 (+0.2) 95.9 (+2.1)
Llama-3.2-1B [MetaLlama3.2_2024] 74.3 75.8 (+1.5) 71.0 (-3.3) 80.9 (+6.6)
Llama-3.2-3B [MetaLlama3.2_2024] 78.4 78.8 (+0.4) 74.5 (-3.9) 81.1 (+2.7)
@@
USPTO50k GPT-5 [openai2025gpt5] 9.0 14.0 (+5.0) 12.0 (+3.0) 16.0 (+7.0)
Gemini-2.5-Pro [google2025gemini2_5_pro] 9.0 15.0 (+6.0) 12.0 (+3.0) 18.0 (+9.0)
Claude-Sonnet-4.5 [anthropic2025sonnet4.5] 20.0 24.0 (+4.0) 23.0 (+3.0) 30.0 (+10.0)
@@
ProteinGym DeepSeek-V3.1-Terminus [DeepSeekV3.1Terminus] 47.9 48.9 (+1.0) 48.6 (+0.7) 56.8 (+8.9)
GPT-OSS-120B [openai2025gptoss120bgptoss20bmodel] 47.7 49.8 (+2.1) 51.5 (+3.8) 57.3 (+9.6)
Claude-Sonnet-4 [anthropic2025claude4] 46.0 49.8 (+3.8) 50.2 (+4.2) 59.7 (+13.7)
@@
As shown in Table 1 and Figure 1, our FLEX paradigm consistently and significantly outperforms all baselines across both mathematical reasoning benchmarks. On the challenging AIME25 dataset, FLEX boosts the accuracy of Claude-Sonnet-4 from a $40.0\%$ baseline to an impressive $63.3\%$ after learning from only 49 examples. This substantial $58.7\%$ relative improvement demonstrates its powerful learning capability. Similarly, on GSM8k, FLEX consistently enhances the performance of already strong models, such as elevating GPT-4’s accuracy to $95.9\%$. These results demonstrate that our parameter-free paradigm can yield substantial gains in reasoning, an advancement previously achieved primarily through parameter-based learning methods [deepseekai2025deepseekr1incentivizingreasoningcapability, yu2025dapoopensourcellmreinforcement].
@@ 5 Discussion
The aspiration to create intelligent agents that evolve through lived experience is a long-standing and central theme in artificial intelligence [turing2009, Russell2021AIMA]. This discussion aims to connect our research to this enduring quest, reflecting on the potential futures it might suggest and the historical context from which it emerges. We hope to offer a thoughtful perspective on the path ahead, framing our contributions not as conclusions, but as a humble step in an ongoing, collective journey.
@@
A central goal in AI is to enable agents to learn continuously from their interactions, much like living organisms do [sutton2023albertaplanairesearch, wang2024comprehensivesurveycontinuallearning]. This pursuit hints at a future where AI systems are not static artifacts, but dynamic partners capable of adapting in real time. For years, the research community has explored various avenues toward this goal, often seeking learning paradigms that are more "forward-looking" than traditional back-propagation [hinton2022forwardforwardalgorithmpreliminaryinvestigations]. The semantic capabilities of modern LLMs [openai2025gpt5, anthropic2025sonnet4.5, anthropic2025claude4, xai2025grok4], however, offer a new lens for this quest. It becomes possible to explore learning not only as a numerical optimization task but as a process of reasoned reflection. Our explorations in this work suggest this is a worthwhile direction, raising the possibility for a new generation of AI systems whose learning processes are inherently more auditable and aligned with a lifelong learning model.
@@
Beyond individual capability, another grand challenge is fostering a form of collective intelligence, mirroring how human progress is built upon shared knowledge [10.1145/3368986]. One can envision a future where a global network of specialized agents collaborate, forming a collective scientific mind to accelerate progress on humanity’s greatest challenges. A key historical obstacle to this vision has been the absence of a robust medium to inherit wisdom across distinct AI agents [CAMPBELL200257, silver2017masteringchessshogiselfplay]. The concept of decoupling learned experience from an agent’s internal parameters represents a hopeful path forward. Our explorations in this direction (Fig 6) suggest that distilled experience can indeed serve as a viable medium for inheritance, adding a small contribution to the larger effort of building a foundation upon which a more synergistic AI ecosystem might one day be built.
@@
Integral to the future of AI is the development of systems whose reasoning is transparent. This points toward a future where human-AI collaboration is built on a foundation of shared understanding, which is crucial for their integration into critical societal functions. The "black box" nature of many deep learning models has presented a persistent barrier to this goal [https://doi.org/10.1002/ail2.61, gohel2021explainableaicurrentstatus, hsieh2024comprehensiveguideexplainableai]. An experience-driven learning paradigm inherently addresses this challenge. When an agent’s growth is chronicled as a series of explicit, human-readable experiences, its evolution becomes an open book. This transparency provides a direct mechanism for meaningful human-in-the-loop interaction [Wu_2022], allowing experts to understand, guide, and even enrich an agent’s learning journey. It represents a step toward a more white-box model of AI development, where the process of learning is as important as the final performance.
@@ 6.1 Learning Paradigms
Supervised fine-tuning (SFT) adapts LLMs to downstream distributions via gradient-based optimization on labeled data or human-written instuctions [howard2018universal, devlin2019bert, radford2019language, raffel2020exploring]. It originates from the pretrain-finetune paradigm established by ULMFiT [howard2018universal], BERT [devlin2019bert] and GPT-2 [radford2019language], and was later formalized as the first stage of alignment training in InstructGPT [ouyang2022training] and FLAN [wei2021finetuned]. SFT further demonstrates remarkable progress across multiple fronts, including instruction following [ouyang2022training, sanh2021multitask, wei2021finetuned], problem solving [toshniwal2024openmathinstruct, ahn2024large, luo2023wizardmath], domain-specific adaptation [luo2022biogpt, zhang2024chemllm, liu2023fingpt, zhou2024lawgpt], and so on [chen2021evaluating, roziere2023code, li2023starcoder, fried2022incoder].
Reinforcement learning (RL) further optimizes LLMs through policy gradient [schulman2017proximal, sutton1999policy, schulman2015trust], aligning model behaviors with human preferences or task-specific rewards, as exemplified by RLHF [christiano2017deep, ziegler2019fine, stiennon2020learning, ouyang2022training, rafailov2023direct] and RLAIF [bai2022training, lee2023rlaif] Beyond alignment, RL has recently proven effective in enhancing model reasoning capabilities on challenging tasks, such as mathematical reasoning [shao2024deepseekmath, guo2025deepseek, yu2025dapo, yue2025vapo], code generation [le2022coderl, shojaee2023execution, wang2024rlcoder], as well as broader structured decision-making tasks [cui2025entropy, hu2025open, zheng2025group, zhang2025survey].
Despite the effectiveness, gradient-based learning paradigms suffer from heavy computational cost and catastrophic forgetting [goodfellow2013empirical, hadi2023large], and the resulting models remain static after training, unable to continually evolve through interaction with the environment [abbas2023loss].
Non-parametric adaptation, including prompt engineering [sahoo2024systematic] and in-context learning [dong2022survey], is another paradigm without gradient-based optimization. Prompt engineering focuses on explicitly designing task instructions or linguistic cues that steer model outputs through semantic conditioning, eliciting step-by-step reasoning without parameter updates [kojima2022large, reynolds2021prompt, wang2022self, yao2023tree, besta2024graph, sahoo2024systematic]. In-context learning adapts frozen LLMs by conditioning them on exemplars of input–output pairs or intermediate reasoning trajectories [brown2020language, wei2022chain, dong2022survey]. However, these works focus on optimizing prompt or context, failing to make model learn from the experiences through interaction with environment.

### v2

@@ 1 Introduction
Autonomous agents have achieved remarkable advances across diverse domains, including code generation (e.g., Claude Code [1]), scientific discovery (e.g., Biomni [2], STELLA [3], and SciMaster [4]), and in-depth research (e.g., Tongyi, OpenAI, and Gemini DeepResearch [5, 6, 7]), showcasing their immense potential to solve complex, open-ended real-world tasks end to end.
However, pretrained agents remain static with frozen parameters, fundamentally precludes learning from on-the-fly experience concluded during trial-and-error, causing substantial performance degradation on challenging or untrained tasks [8, 9]. While gradient-based learning has underpinned model optimization [10, 11, 12], it is inherently ill-suited for continuous agent evolution due to three key obstacles: (i) back-propagation incurs prohibitive computational costs [13, 14]; (ii) parameter-centric paradigm suffers from catastrophic forgetting [15, 16], failing to effectively learn form accumulated experiences over times; (iii) the closed-source nature of most state-of-the-art LLMs [17, 18, 19, 20, 21] renders direct parameter optimization infeasible.
Recent self-evolving paradigms attempt to overcome this limitation by evolving non-parametric components (i.e., prompts [22, 23], workflows [24, 25, 26], and tools [27, 28]) through environmental feedback and textual gradient [22, 29, 30]. Nevertheless, these approaches encounter three vital bottlenecks towards continuous agent evolution: (i) their evolving components are task-specific, thus they cannot generalize across tasks, limiting cross-task evolution. (ii) Their evolving components (i.e. prompts, workflows, and tools) are not continuously scalable, thus they can only leverage a limited set of experiences, preventing performance from scaling with accumulated knowledge. (iii) Their evolving components are model-specific, thus new agents must interact from scratch during deployment, unable to continuously evolve from agents’ previous off-policy experiences, leading to substantial computational overhead.
@@
To validate the general effectiveness of FLEX paradigm, we conduct extensive experiments across diverse challenging scientific domains, including Olympiad-level mathematics (AIME25 [31]), chemical retrosynthesis (USPTO50k [32]), and protein fitness prediction (ProteinGym [33]). FLEX demonstrates substantial and consistent improvements on these tasks, from $40\%$ to $63\%$ on AIME25, $20\%$ to $30\%$ on USPTO50k, and $46\%$ to $60\%$ on ProteinGym, exhibiting great enhancement in the capacity of reasoning and knowledge leverage.
@@ 3.1 Extensive Experience Exploration
To enable extensive high-quality exploration, we design a dual-scaling mechanism that combines parallel and sequential scaling. For parallel scaling, multiple reasoning trajectories are sampled in parallel for the same input through rejection sampling [34, 35, 36], allowing diverse reasoning hypotheses to be explored and high-quality outputs to be retained. For sequential scaling, the actor iteratively improves upon its previous output under the critic’s semantic feedback [22, 29], where the critic provides constructive guidance and refinement cues derived from the prior round. This cooperative exploration strategy balances breadth and depth, ensuring both coverage of alternative reasoning paths and progressive improvement over iterations.
@@
To ensure that collected experiences contribute meaningfully to learning, we design an iterative refinement mechanism driven by semantic evaluation. For reasoning trajectories that lead to incorrect conclusions, the critic first identifies the underlying causes of failure and summarizes them into abstract, task-agnostic improvement suggestions. These generalized feedback signals, denoted as new experiences $E_{t}$, are fed back into the actor together with its previous reasoning trace to initiate the next actor–critic iteration. This process continues until the actor produces a correct or sufficiently improved reasoning trajectory, or a predefined iteration limit is reached. In this loop, the critic’s feedback functions as a semantic update signal [22, 29, 37] that guides the actor’s reasoning refinement, analogous to the gradient signal in traditional optimization, yet expressed in natural language semantics rather than numerical differentials. Unlike backpropagation-based learning, which relies on scalar gradients with limited interpretive bandwidth and heavy computational budget, this forward semantic feedback mechanism provides rich interpretable guidance without costly parameter updating, aligning with the cognitive characteristics of human-like learning from experience.
@@ 3.2 Experience Library Evolution
During inference, the experience library is encapsulated as an interactive tool callable by the agent [38, 28]. Instead of static vector-based semantic-similarity search [39, 40, 41], the agent performs contextualized retrieval conditioned on the specific query and its current reasoning state. This allows the agent to interpret relevance not only by semantic similarity but also by contextual consistency and goal alignment, thereby avoiding the pitfalls of retrieving semantically similar yet pragmatically contradictory entries. Retrieval proceeds hierarchically, the agent first identifies relevant high-level strategies, then locates corresponding procedural patterns, and finally accesses concrete case examples within each pattern. At each stage, the top-$k$ (typically $k=5$) most relevant entries are selected to balance precision and efficiency. The retrieved experiences support the agent’s ongoing reasoning, reflective self-evaluation, and corrective refinement. Moreover, retrieval is not restricted to pre-inference access, instead, the agent can dynamically query the library during reasoning, enabling adaptive integration of prior knowledge into evolving cognitive trajectories.
@@ 4.1 Experimental Setup
We evaluate the proposed FLEX paradigm across four challenging benchmarks spanning three scientific domains (i.e., mathematics, chemistry, and biology) to assess both reasoning capability and domain adaptability. Specifically: (1) AIME25 competition dataset [31] for complicated mathematical reasoning; (2) GSM8k dataset [42] multi-step arithmetic reasoning; (3) USPTO50k [32] benchmark for canonical single-step retrosynthesis; (4) ProteinGym [33] benchmark for protein fitness prediction.
@@
Our evaluation spans a range of base models, tested on four distinct experimental configurations for comparison: (1) Vanilla LLM, (2) LLM with In-Context Learning (ICL) [43], (3) LLM agent with reasoning-and-acting workflow (Agent) [44], and (4) our proposed FLEX. For FLEX, each model constructs an experience library during training, which is later utilized at test-time. The library structure is adapted to the benchmark: hierarchical for AIME25 and USPTO50k, and non-hierarchical for GSM8k and ProteinGym.
@@
For AIME25, the training was conducted on 49 historical problems from AIME83 to AIME24. On GSM8k, we adhere to the official training and testing splits [42]. For USPTO50k, 100 instances are randomly sampled from the original 5k test split for evaluation, and 50 instances from the training split are randomly sampled for training. For ProteinGym, 100 mutational sequences are sampled for each wild-type protein target as training set, which is on average 1.47% of the available data. This design reflects practical constraints in protein engineering, where only a limited number of labeled mutations are known, while the vast majority of sequence space remains unexplored.
@@ 4.2 Main Results
AIME25 Claude-Sonnet-4 [20] 40.0 30.0 (-10.0) 50.0 (+10.0) 63.3 (+23.3)
DeepSeek-V3.1-Terminus [45] 56.7 53.3 (-3.3) 60.0 (+3.3) 66.6 (+10.0)
GSM8k GPT-3.5 [46] 80.8 81.4 (+0.6) 78.5 (-2.3) 83.3 (+3.3)
GPT-4 [47] 93.8 94.2 (+0.4) 94.0 (+0.2) 95.9 (+2.1)
Llama-3.2-1B [48] 74.3 75.8 (+1.5) 71.0 (-3.3) 80.9 (+6.6)
Llama-3.2-3B [48] 78.4 78.8 (+0.4) 74.5 (-3.9) 81.1 (+2.7)
@@
USPTO50k GPT-5 [18] 9.0 14.0 (+5.0) 12.0 (+3.0) 16.0 (+7.0)
Gemini-2.5-Pro [17] 9.0 15.0 (+6.0) 12.0 (+3.0) 18.0 (+9.0)
Claude-Sonnet-4.5 [19] 20.0 24.0 (+4.0) 23.0 (+3.0) 30.0 (+10.0)
@@
ProteinGym DeepSeek-V3.1-Terminus [45] 47.9 48.9 (+1.0) 48.6 (+0.7) 56.8 (+8.9)
GPT-OSS-120B [49] 47.7 49.8 (+2.1) 51.5 (+3.8) 57.3 (+9.6)
Claude-Sonnet-4 [20] 46.0 49.8 (+3.8) 50.2 (+4.2) 59.7 (+13.7)
@@
As shown in Table 1 and Figure 1, our FLEX paradigm consistently and significantly outperforms all baselines across both mathematical reasoning benchmarks. On the challenging AIME25 dataset, FLEX boosts the accuracy of Claude-Sonnet-4 from a $40.0\%$ baseline to an impressive $63.3\%$ after learning from only 49 examples. This substantial $58.7\%$ relative improvement demonstrates its powerful learning capability. Similarly, on GSM8k, FLEX consistently enhances the performance of already strong models, such as elevating GPT-4’s accuracy to $95.9\%$. These results demonstrate that our parameter-free paradigm can yield substantial gains in reasoning, an advancement previously achieved primarily through parameter-based learning methods [50, 51].
@@ 5 Discussion
The aspiration to create intelligent agents that evolve through lived experience is a long-standing and central theme in artificial intelligence [52, 53]. This discussion aims to connect our research to this enduring quest, reflecting on the potential futures it might suggest and the historical context from which it emerges. We hope to offer a thoughtful perspective on the path ahead, framing our contributions not as conclusions, but as a humble step in an ongoing, collective journey.
@@
A central goal in AI is to enable agents to learn continuously from their interactions, much like living organisms do [54, 55]. This pursuit hints at a future where AI systems are not static artifacts, but dynamic partners capable of adapting in real time. For years, the research community has explored various avenues toward this goal, often seeking learning paradigms that are more "forward-looking" than traditional back-propagation [56]. The semantic capabilities of modern LLMs [18, 19, 20, 21], however, offer a new lens for this quest. It becomes possible to explore learning not only as a numerical optimization task but as a process of reasoned reflection. Our explorations in this work suggest this is a worthwhile direction, raising the possibility for a new generation of AI systems whose learning processes are inherently more auditable and aligned with a lifelong learning model.
@@
Beyond individual capability, another grand challenge is fostering a form of collective intelligence, mirroring how human progress is built upon shared knowledge [57]. One can envision a future where a global network of specialized agents collaborate, forming a collective scientific mind to accelerate progress on humanity’s greatest challenges. A key historical obstacle to this vision has been the absence of a robust medium to inherit wisdom across distinct AI agents [58, 59]. The concept of decoupling learned experience from an agent’s internal parameters represents a hopeful path forward. Our explorations in this direction (Fig 6) suggest that distilled experience can indeed serve as a viable medium for inheritance, adding a small contribution to the larger effort of building a foundation upon which a more synergistic AI ecosystem might one day be built.
@@
Integral to the future of AI is the development of systems whose reasoning is transparent. This points toward a future where human-AI collaboration is built on a foundation of shared understanding, which is crucial for their integration into critical societal functions. The "black box" nature of many deep learning models has presented a persistent barrier to this goal [60, 61, 62]. An experience-driven learning paradigm inherently addresses this challenge. When an agent’s growth is chronicled as a series of explicit, human-readable experiences, its evolution becomes an open book. This transparency provides a direct mechanism for meaningful human-in-the-loop interaction [63], allowing experts to understand, guide, and even enrich an agent’s learning journey. It represents a step toward a more white-box model of AI development, where the process of learning is as important as the final performance.
@@ 6.1 Learning Paradigms
Supervised fine-tuning (SFT) adapts LLMs to downstream distributions via gradient-based optimization on labeled data or human-written instuctions [64, 65, 66, 67]. It originates from the pretrain-finetune paradigm established by ULMFiT [64], BERT [65] and GPT-2 [66], and was later formalized as the first stage of alignment training in InstructGPT [68] and FLAN [69]. SFT further demonstrates remarkable progress across multiple fronts, including instruction following [68, 70, 69], problem solving [71, 72, 73], domain-specific adaptation [74, 75, 76, 77], and so on [78, 79, 80, 81].
Reinforcement learning (RL) further optimizes LLMs through policy gradient [82, 83, 84], aligning model behaviors with human preferences or task-specific rewards, as exemplified by RLHF [85, 86, 87, 68, 88] and RLAIF [89, 90] Beyond alignment, RL has recently proven effective in enhancing model reasoning capabilities on challenging tasks, such as mathematical reasoning [91, 92, 93, 94], code generation [95, 96, 97], as well as broader structured decision-making tasks [98, 99, 100, 101].
Despite the effectiveness, gradient-based learning paradigms suffer from heavy computational cost and catastrophic forgetting [13, 16], and the resulting models remain static after training, unable to continually evolve through interaction with the environment [15].
Non-parametric adaptation, including prompt engineering [102] and in-context learning [43], is another paradigm without gradient-based optimization. Prompt engineering focuses on explicitly designing task instructions or linguistic cues that steer model outputs through semantic conditioning, eliciting step-by-step reasoning without parameter updates [103, 104, 105, 106, 107, 102]. In-context learning adapts frozen LLMs by conditioning them on exemplars of input–output pairs or intermediate reasoning trajectories [108, 109, 43]. However, these works focus on optimizing prompt or context, failing to make model learn from the experiences through interaction with environment.